A collection of 33 coins, consisting of nickels, dimes, and quarters, has a value of $3.30. If there are three times as many nickels as quarters, and one-half as many dimes as nickels, how many coins of each kind are there.
Answer by mananth(16946)

A collection of 33 coins, consisting of nickels, dimes, and quarters, has a value of $3.30. If there are three times as many nickels as quarters, and one-half as many dimes as nickels, how many coins of each kind are there.
Quarters =x
Nickels = 3x
dimes = 3x/2
$3.30 = 330 cents
25x+5*3x+10*3x/2=330
25x+15x+15x=330
55x=330
/55
x=6 quarters
Nickels = 3x=>18
Dimes= 3x/2=>9